# bibliography.bib

### v1

@@ Abstract
Autonomous vehicles (AV) offer a cost-effective solution for scientific missions such as underwater tracking. Recently, reinforcement learning (RL) has emerged as a powerful method for controlling AVs in complex marine environments. However, scaling these techniques to a fleet—essential for multi-target tracking or targets with rapid, unpredictable motion—presents significant computational challenges. Multi-Agent Reinforcement Learning (MARL) is notoriously sample-inefficient, and while high-fidelity simulators like Gazebo’s LRAUV provide 100× faster-than-real-time single-robot simulations, they offer no significant speedup for multi-vehicle scenarios, making MARL training impractical. To address these limitations, we propose an iterative distillation method that transfers high-fidelity simulations into a simplified, GPU-accelerated environment while preserving high-level dynamics. This approach achieves up to a 30,000× speedup over Gazebo through parallelization, enabling efficient training via end-to-end GPU acceleration. Additionally, we introduce a novel Transformer-based architecture (TransfMAPPO) that learns multi-agent policies invariant to the number of agents and targets, significantly improving sample efficiency. Following large-scale curriculum learning conducted entirely on GPU, we perform extensive evaluations in Gazebo, demonstrating that our method maintains tracking errors below 5 meters over extended durations, even in the presence of multiple fast-moving targets. This work bridges the gap between large-scale MARL training and high-fidelity deployment, providing a scalable framework for autonomous fleet control in real-world sea missions.
@@ I INTRODUCTION
Figure 1: Overview of our Training and Evaluation Pipeline.
@@
Underwater tracking (UT) is essential to advance marine research missions, such as tracking marine species and oceanographic phenomena [zarokanellos2022frontal, masmitja2020mobile], monitoring large-scale data collection using autonomous underwater vehicles (AUV) [Yanwu2021], and managing marine protected areas [vigo2021spatial]. However, underwater communication systems face challenges such as low reliability and bandwidth, as GPS is ineffective [heidemann2012underwater]. Acoustic tracking using AUVs or autonomous surface vehicles (ASVs) offers a promising solution, improving efficiency and reducing costs compared to traditional fixed equipment [masmitja2020mobile, Yanwu2021]. Yet, challenges like unreliable communication, complex environment dynamics, and energy constraints necessitate advanced motion planning to enhance AVs’ tracking capabilities.
Reinforcement learning (RL) enables AVs to learn optimal navigation planning through trial and error, offering dynamic responses to environmental conditions compared to pre-programmed approaches [cote2019characterizing]. Recent work by [ivanscience] demonstrated the potential of RL by training a Soft Actor-Critic (SAC) agent in a simulator to control an ASV tracking an underwater target using range-only acoustic data. While successful, the model was limited to a single target moving linearly at modest speeds (up to 0.3 times the agent’s velocity).
Scaling this approach to more complex scenarios requires multi-agent systems. Multi-Agent Reinforcement Learning (MARL) could enable learning an optimal cooperative tracking policy by sharing tracking data collected simultaneously from different locations. However, deploying such a complex system in the sea requires preliminary extensive testing. Simulators like Gazebo’s Long-Range AUV simulator (LRAUV Sim) [lrauv_sim_icra], which models hydrodynamics, acoustic communication, and marine sensors, are crucial tools in this phase. Unfortunately, LRAUV Sim offers only a  10x faster-than-real-time speedup in multi-robot simulations, and its C++-based control complicates integration with RL methods, meaning that MARL cannot be applied directly in such a simulator.
@@
  • PyLrauv  [^1] https://github.com/mttga/pylrauv: An open-source Python package using ROS2 [ros2] to control LRAUV Sim with a Gym-like interface. PyLrauv enables seamless control and observation of a variable number of agents and targets while ensuring high-fidelity simulation via Gazebo. It also supports direct deployment in real missions.
  • Environment Distillation: A method to create a simplified, GPU-accelerated environment using JAX, achieving up to 30,000x speedup over Gazebo and allowing to train effective policies in minutes. This distilled environment, which we integrate into the popular MARL library JaxMARL [flair2023jaxmarl]  [^2] https://github.com/FLAIROx/JaxMARL, allows rapid training while maintaining high-level equivalence to PyLrauv.
  • TransfMAPPO: A novel method integrates Transformers into MAPPO [mappo] to train policies independent of the number of entities in the environment. TransfMAPPO enables a curriculum learning (CL) pipeline that learns robust policies in complex multi-agent scenarios, where traditional from-scratch methods like MAPPO are ineffective.
  • High-Fidelity Evaluation: We demonstrate our pipeline effectiveness by training TransfMAPPO purely on GPU on our distilled environment and testing it in high-fidelity simulations using PyLrauv: our trained models can track erratically moving targets at 0.8 times their speed and multiple targets simultaneously. We open-source TransfMAPPO weights for use in LRAUV Sim and real missions via PyLrauv.
@@ II RELATED WORK
Early research in UT with autonomous vehicles has relied on traditional control strategies for AVs [heidemann2012underwater, almeida2017], with more recent work incorporating reinforcement learning (RL) to improve navigation and tracking robustness [ivanscience]. However, these studies have mainly focused on single-agent scenarios or environments with limited target dynamics.
MARL has emerged as a natural extension of RL for learning cooperative policies, yet its sample inefficiency and scalability remain challenging. Transformer-based methods, such as TransfQMix [transfqmix], decouple policy learning from the number of entities by leveraging specialized transformer modules, allowing, in principle, the training of the same network in progressively more challenging multi-agent scenarios. Our TransfMAPPO architecture is inspired by TransfQMix, but uses MAPPO instead of DQN [dqn2016] as the RL backbone due to its natural compatibility with parallelized environments. Moreover, TransfMAPPO does not use a centralized hypernetwork but rather a centralized critic, simplifying the model structure. Finally, TransfMAPPO is also applicable to continuous action spaces, while TransfQMix is limited to discrete.
Frameworks like JaxMARL [flair2023jaxmarl] have demonstrated the benefits of GPU-accelerated MARL, albeit in relatively simple environments, while packages such as IsaacGym [isaacgym] provide more advanced GPU-based simulation of real robots but do not offer the hydrodynamic models and communication protocols necessary to bridge the sim-to-real gap.
Our pipeline can be associated with model-based RL [model_based_1, model_based_2, model_based_3], which aim to learn a world model to improve and speed up RL training. Our method creates a simplified, GPU-accelerated physical model from a high-fidelity simulator, but instead of focusing on step-wise alignment between the world model and the real world, it focuses on preserving high-level dynamics at a different time scale with respect to a more complex simulator. Moreover, our distilled environment is not accessible by our learning algorithm, as is usually the case in model-based RL; instead, we simply use model-free RL on top of it.
@@ III BACKGROUND
III-A Underwater Tracking with Autonomous Vehicles
Underwater acoustic tracking with autonomous vehicles is a member of the family of Range Only Single Beacon (ROSB) localization techniques [rosb, av_range_only], which use range-only acoustic observations gathered over time to locate a target. For static targets, the trilateration problem can be linearised by least squares (LS) algorithms, which assume static conditions and incur errors when targets move [ThesisLisboa]. For dynamic scenarios, particle filters (PF) use Bayesian estimation, representing potential target states with weighted particles updated via motion models and resampled based on measurement likelihoods [ROSB1, ROSB2].
@@ III-B Multi-Agent Reinforcement Learning
Cooperative multi-agent tasks are formalized as decentralized partially observable Markov decision processes (Dec-POMDP) [dec-Pompd], defined by the tuple $G=\langle S,\mathbf{U},P,r,Z,O,n,\gamma\rangle$, where $S$ is the global state space, $\mathbf{U}=U^{n}$ is the joint action space for $n$ agents (each selecting actions $u^{a}\in U$), $P(s^{\prime}\mid s,\mathbf{u})$ is the state transition function, $r(s,\mathbf{u})$ is the shared reward function, $Z$ is the observation space for each agent, $O(s,a)$ defines the observation $z^{a}\in Z$ for agent $a$, and $\gamma\in[0,1)$ is the discount factor. At each timestep, agent $a$ receives a partial observation $z^{a}=O(s,a)$ and maintains an action-observation history $\tau^{a}=(z_{0}^{a},u_{0}^{a},\dots,z_{t}^{a})\in\mathcal{T}$. The agent’s policy $\pi^{a}(u^{a}\mid\tau^{a}):\mathcal{T}\times U\rightarrow[0,1]$ maps this history to a distribution over actions. The agents collectively aim to maximize the expected discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. A key paradigm in MARL is centralized training with decentralized execution (CTDE) [rashid2018qmixmonotonicvaluefunction], where agents leverage centralized information during training but act independently during execution. This allows the use of a centralized critic that estimates the value function $V(s)$ using the global state $s$, while policies $\pi^{a}$ remain decentralized, depending only on local histories $\tau^{a}$.
MAPPO [mappo] extends Proximal Policy Optimization (PPO) [ppo] to multi-agent settings. Each agent’s policy (actor) $\pi_{\theta}^{a}$ and centralized value function (critic) $V_{\phi}$ are optimized jointly. The actor is updated using the clipped surrogate objective:
@@
where $\rho_{t}(\theta)=\frac{\pi_{\theta}(u_{t}\mid\tau_{t})}{\pi_{\theta_{\text{old}}}(u_{t}\mid\tau_{t})}$ is the probability ratio, $\hat{A}_{t}$ is the advantage estimate, and $\epsilon$ is a clipping hyperparameter. The critic minimizes the temporal difference error:
$\mathcal{L}^{VF}(\phi)=\mathbb{E}_{t}\left[\left(V_{\phi}(s_{t})-\hat{R}_{t}\right)^{2}\right],$
where $\hat{R}_{t}$ is the discounted return. During execution, only the actors are used, enabling decentralized control.
@@ IV-A PyLrauv
PyLrauv provides a Python interface to LRAUV Sim by leveraging ROS 2 [ros2]. This is achieved by integrating a set of additional communication handlers into the official Gazebo C++ library, ros_gz [^3] https://github.com/gazebosim/ros_gz. The integration allows control of the LRAUV robots with python using rclpy  [^4] https://github.com/ros2/rclpy. We have developed a set of Python controllers to send commands to the LRAUV vehicles in Gazebo and observe their published states. Since the actual LRAUV vehicles use the same technology, the controllers provided within PyLrauv can also be used to control real LRAUV robots. Controlling multiple vehicles simulated in Gazebo in a Gym-like fashion is straightforward, as demonstrated by the following code:
@@
IV-B UTracking Environment
UTrackingEnv is a Gym-like environment managing agent and target commands, communication, tracking, and RL tasks like observation building and reward calculations. Agents communicate via the Gazebo LRAUV Acoustic Comms Plugin and collect target ranges using the LRAUV Range Bearing Plugin. At the start of an episode, agents are spawned on the sea surface and targets at random depths, ensuring a minimum distance between each entity of 50 meters and maximum of 200. Both agents and targets are LRAUV vehicles. Each step lasts 30 seconds in our experiments, at the beginning of which agents listen for range signals from targets, and at the end broadcast their locations and observations to other agents. Each agent maintain its own tracking model (LS or PF) for each target, updated with all the information received. This ensures a genuine decentralization of the system. Targets are tracked in 2D space, as depth is typically known [ivanscience].
@@ IV-B UTracking Environment
The agents receive their partial observation of the environment at the end of the step phase, which includes the relative distances to all other entities in the 3D space. For other agents, these distances are calculated using the information they have received from them (if available), and for targets, the distances are calculated from the tracking predictions generated by their own tracking modules. [ivanscience] incorporates the range observations into observation vectors, which may be helpful, but may also cause agents to overfit to training dynamics and overlook tracking information. To prevent this, we limit observations to tracking-derived values, ensuring agents rely on tracking data and its inherent errors. The environment also returns the global state of the environment, which includes the true 3D position of each entity, its velocity, and its direction.
@@
[…]
$r=\frac{1}{N}\sum_{i=1}^{N}r_{i},$ (3)
where $N$ is the number of targets to track, i.e. the reward is normalized by the number of targets to be always in the interval $[0,1]$, so that learned value functions can be transferred across scenarios with different numbers of targets. In addition, a crash penalty is always applied if the distance between any two agents falls below a minimum valid distance, $d_{safe}$, i.e., $r=-1\quad\text{if}\quad\min_{i\neq j}\{d_{ij}\}<d_{safe}.$
@@
To simplify the action space of our multi-agent system, we keep the agents’ velocity constant and allow them to control only the rudder. The rudder is discretized into five values corresponding to equidistant angles ranging from $-0.24$ to $0.24$ radians. Agents can adjust the rudder only to the next closest two angles, preventing abrupt changes in direction.
@@
The episode ends when the maximum number of steps is reached. The environment is not truncated when crashes occur or targets are lost, as we experimentally observed that truncating episodes would create training instabilities.
Why UTracking is a difficult environment?
  • Partial observability: Agents detect targets only within 450 m via noisy acoustic signals and rely on intermittent and unstable communication (1500 m range) for shared tracking data.
  • Stochasticity: Sensor noise, communication dropouts, environmental perturbations, and unpredictable target motions amplify uncertainty.
  • Coordination challenge: Agents must coordinate almost perfectly especially when tracking multiple rapid-moving targets, since a lack of coordination could result in rapidly loosing a target.
  • Long horizons: Hour-scale missions (corresponding to thousands of steps) require learning delayed consequences of actions, which can be difficult in Multi-Agent scenarios given their non-stationarity. Moreover, they need to be robust to error accumulation in tracking models.
IV-C Environment Distillation: JaxLrauv UTracking

### v2

# Scaling Multi-Agent Reinforcement Learning for Underwater Acoustic Tracking via Autonomous Vehicles

@@ Abstract
Autonomous vehicles (AVs) offer a cost-effective solution for scientific missions such as underwater tracking. Reinforcement learning (RL) has emerged as a powerful method for controlling AVs, but scaling to fleets (essential for multi-target tracking or rapidly moving targets) is challenging. Multi-Agent RL (MARL) is notoriously sample-inefficient, and while high-fidelity simulators like Gazebo’s LRAUV provide up to 100× faster-than-real-time single-robot simulations, they offer little speedup in multi-vehicle scenarios, making MARL training impractical. Yet, high-fidelity simulation is crucial to test complex policies and close the sim-to-real gap. To address these limitations, we develop a GPU-accelerated environment that achieves up to 30,000× speedup over Gazebo while preserving its dynamics. This enables fast, end-to-end GPU training and seamless transfer to Gazebo for evaluation. We also introduce a Transformer-based architecture (TransfMAPPO) that learns policies invariant to fleet size and number of targets, enabling curriculum learning to train larger fleets on increasingly complex scenarios. After large-scale GPU training, we perform extensive evaluations in Gazebo, showing our method maintains tracking errors below 5m even with multiple fast-moving targets.
@@ I INTRODUCTION
Underwater tracking (UT) is essential to advance marine research missions, such as tracking marine species and oceanographic phenomena [1, 2], monitoring large-scale data collection using autonomous underwater vehicles (AUV) [3], and managing marine protected areas [4]. However, underwater communication systems face challenges such as low reliability and bandwidth, as GPS is ineffective [5]. Acoustic tracking using AUVs or autonomous surface vehicles (ASVs) offers a promising solution, improving efficiency and reducing costs compared to traditional fixed equipment [2, 3]. Yet, challenges like unreliable communication, complex environment dynamics, and energy constraints necessitate advanced motion planning to enhance AVs’ tracking capabilities.
Figure 1: Overview of our training and evaluation pipeline. PyLrauv $(a)$ is a new Python package to control multiple robots in the C++ high-fidelity LRAUV simulator. JaxLrauv $(b)$ is a GPU-accelerated, simplified environment that supports massive parallelization while preserving the dynamics of the LRAUV simulator. TransfMAPPO $(c)$ employs transformers to train progressively larger fleets of vehicles to coordinate via curriculum learning. The final policies trained in JaxLrauv with TransfMAPPO are then evaluated in the realistic LRAUV simulator prior to real-world deployment.
@@
Reinforcement learning (RL) enables AVs to learn optimal navigation strategies through trial and error, providing dynamic responses to environmental conditions compared to pre-programmed approaches [6]. Preliminary work by [7] demonstrated the potential of using RL to train an ASV policy for tracking an underwater target using range-only acoustic data. More recent studies [8, 9] have shown that Multi-Agent Reinforcement Learning (MARL) can enable cooperative tracking by sharing data collected simultaneously from different locations. However, these MARL studies were conducted in very simple, abstract environments. Deploying such a complex system in the sea requires extensive preliminary testing. Simulators such as Gazebo’s Long-Range AUV simulator (LRAUV Sim) [10], which models hydrodynamics, acoustic communication, and marine sensors of real vehicles, are crucial tools for this phase. Unfortunately, LRAUV Sim provides only about a $10\times$ faster-than-real-time speedup in multi-robot simulations, and its C++-based control complicates integration with RL methods, making direct application of MARL impractical in this simulator.
@@
[…]
  • Agent-Target Invariant Policies with Transformers and Curriculum Learning: By integrating Transformers into a novel variant of MAPPO [13] (TransfMAPPO), we train policies that are invariant to the number of agents and targets. This enables a curriculum learning (CL) pipeline that progressively trains larger fleets in increasingly complex multi-target scenarios. TransfMAPPO learns robust cooperative policies where traditional from-scratch approaches such as MAPPO fail. Thanks to the combination of fast training and curriculum learning, we obtain a general policy capable of tracking up to 5 simultaneous targets with only 5 vehicles (f.i. Figure 2), whereas state-of-the-art require up to 12 vehicles to track 4 targets.
@@ II RELATED WORK
Early research in UT with autonomous vehicles has relied on traditional control strategies for AVs [5, 14]. A promising new direction was opened by [7], who incorporated RL to improve navigation and tracking robustness, though their work was limited to single vehicles tracking fixed or slowly moving targets. More recent studies [9, 8] have explored the use MARL to train cooperative policies for tracking moving targets, but these methods were implemented in highly abstract, simplified environments. Our proposed pipeline, in contrast, ensures that learned policies are robust to realistic scenarios. Moreover  [9] only demonstrated a cooperative policy with two vehicles tracking a single target. Similarly to us, [8] employed an attention mechanism, but did not train team-size invariant policies. Instead, they trained fixed configurations such as 4 agents tracking 2 targets, 6 agents and 3 targets, or 12 agents and 4 targets. By comparison, our method is able to train policies that generalize to team sizes and target numbers, and we demonstrate robust tracking of 5 targets using only 5 vehicles. Neither of these works provide wall-clock training times or official code-bases to obtain them, but they certainly lack environment vectorization and pure-GPU training, allowing us to safely assert that in a comparison our pipeline would be significantly faster and more robust.
Our TransfMAPPO architecture is inspired by TransfQMix [15], but adopts PPO instead of Q-Learning as the RL backbone due to its natural compatibility with parallelized environments. Moreover, TransfMAPPO replaces the centralized hypernetwork with a centralized critic, simplifying the model structure. Unlike TransfQMix, TransfMAPPO also supports continuous action spaces, rather than being restricted to discrete ones. Several other methods have explored scaling coordination in MARL using attention mechanisms, Graph Neural Networks (GNNs), or Transformers [16, 17, 18, 19], but a detailed comparison with these approaches is beyond the scope of this work. The most similar work to ours that combines transformers with MAPPO is Multi-Agent Transformer [20]. However, in that work transformers are used to model the multi-agent problem as a sequential decision process, whereas in our case we employ them to improve simultaneous coordination. Related to our approach, [21, 22] employ curriculum learning to progressively train policies that adapt to larger team sizes, but these methods have not been evaluated in complex robotic scenarios.
@@ III BACKGROUND
III-A Underwater Localization with Autonomous Vehicles
Underwater acoustic tracking with autonomous vehicles is a member of the family of Range Only Single Beacon (ROSB) localization techniques [23, 24], which use range-only acoustic observations gathered over time to locate a target. For static targets, the trilateration problem can be linearised by least squares (LS) algorithms, which assume static conditions and incur errors when targets move [25]. For dynamic scenarios, particle filters (PF) use Bayesian estimation, representing potential target states with weighted particles updated via motion models and resampled based on measurement likelihoods [26, 27].
@@ III-B Multi-Agent Reinforcement Learning
Cooperative multi-agent tasks are formalized as decentralized partially observable Markov decision processes (Dec-POMDP) [28], defined by the tuple $G=\langle S,\mathbf{U},P,r,Z,O,n,\gamma\rangle$, where $S$ is the global state space, $\mathbf{U}=U^{n}$ is the joint action space for $n$ agents (each selecting actions $u^{a}\in U$), $P(s^{\prime}\mid s,\mathbf{u})$ is the state transition function, $r(s,\mathbf{u})$ is the shared reward function, $Z$ is the observation space for each agent, $O(s,a)$ defines the observation $z^{a}\in Z$ for agent $a$, and $\gamma\in[0,1)$ is the discount factor. At each timestep, agent $a$ receives a partial observation $z^{a}=O(s,a)$ and maintains an action-observation history $\tau^{a}=(z_{0}^{a},u_{0}^{a},\dots,z_{t}^{a})\in\mathcal{T}$. The agent’s policy $\pi^{a}(u^{a}\mid\tau^{a}):\mathcal{T}\times U\rightarrow[0,1]$ maps this history to a distribution over actions. The agents collectively aim to maximize the expected discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. A key paradigm in MARL is centralized training with decentralized execution (CTDE) [29], where agents leverage centralized information during training but act independently during execution. This allows the use of a centralized critic that estimates the value function $V(S)$ using the global state $S$, while policies $\pi^{a}$ remain decentralized, depending only on local histories $\tau^{a}$.
MAPPO [13] extends Proximal Policy Optimization (PPO) [30] to multi-agent settings. Each agent’s policy (actor) $\pi_{\theta}^{a}$ and centralized value function (critic) $V_{\phi}$ are optimized jointly. The actor is updated using the clipped surrogate objective:
@@
where $\rho_{t}(\theta)=\frac{\pi_{\theta}(u_{t}\mid\tau_{t})}{\pi_{\theta_{\text{old}}}(u_{t}\mid\tau_{t})}$ is the probability ratio, $\hat{A}_{t}$ is the advantage estimate, and $\epsilon$ is a clipping hyperparameter. The critic minimizes the temporal difference error: $\mathcal{L}^{VF}(\phi)=\mathbb{E}_{t}\left[\left(V_{\phi}(s_{t})-\hat{R}_{t}\right)^{2}\right]$, where $\hat{R}_{t}$ is the discounted $\lambda$ return. During execution, only the actors are used, enabling decentralized control.
@@ IV-A PyLrauv
PyLrauv provides a Python interface to LRAUV Sim by leveraging ROS 2 [11]. This is achieved by integrating a set of additional communication handlers into the official Gazebo C++ library, ros_gz [^1] https://github.com/gazebosim/ros_gz. The integration allows control of the LRAUV robots with Python using rclpy  [^2] https://github.com/ros2/rclpy. We have developed a set of Python controllers to send commands to the LRAUV vehicles in Gazebo and observe their published states. Since the actual LRAUV vehicles use the same technology, the controllers provided within PyLrauv can also be used to control real LRAUV robots. Controlling multiple vehicles simulated in Gazebo in a Gym-like fashion is straightforward, as demonstrated by the following code:
@@
IV-B PyLrauv UTracking Environment
UTrackingEnv is a Gym-like environment managing agent and target commands, communication, tracking, and RL tasks like observation building and reward calculations. Agents communicate via the Gazebo LRAUV Acoustic Comms Plugin and collect target ranges using the LRAUV Range Bearing Plugin. At the start of an episode, agents are spawned on the sea surface and targets at random depths, ensuring a minimum distance between each entity of 50 m and maximum of 200 m. Both agents and targets are LRAUV vehicles. Each step lasts 30 seconds in our experiments, at the beginning of which agents listen for range signals from targets, and at the end broadcast their locations and observations to other agents (communication phase). Each agent maintains its own tracking model (LS or PF) for each target, updated with all the information received. This ensures a genuine decentralization of the system. Targets are tracked in 2D space, as depth is typically known [7].
@@ IV-B PyLrauv UTracking Environment
The agents receive their partial observation of the environment at the end of the step phase, which includes the relative distances to all other entities in the 3D space. For other agents, these distances are calculated using the information they have received from them (if available), and for targets, the distances are calculated from the tracking predictions generated by their own tracking modules. [7] incorporates the range observations into observation vectors, which may be helpful, but may also cause agents to overfit to training dynamics and overlook tracking information. To prevent this, we limit observations to tracking-derived values, ensuring agents rely on tracking data and its inherent errors. The environment also returns the global state of the environment, which includes the true 3D position of each entity, its velocity, and its direction.
@@
with the threshold set in our experiments to $50\,\text{m}$ (and $100\,\text{m}$ for fast targets). In both cases, the global reward to the system is given by $r=\frac{1}{N}\sum_{i=1}^{N}r_{i},$ where $N$ is the number of targets to track, i.e. the reward is normalized by the number of targets to be always in the interval $[0,1]$, so that learned value functions can be transferred across scenarios with different numbers of targets. In addition, a crash penalty is always applied if the distance between any two agents falls below a minimum valid distance, $d_{safe}$, i.e., $r=-1\quad\text{if}\quad\min_{i\neq j}\{d_{ij}\}<d_{safe}.$
@@
To simplify the action space of our multi-agent system, we keep the agents’ velocity constant and allow them to control only the rudder. The rudder is discretized into five values corresponding to equidistant angles ranging from 0.24-0.24 to $0.24$ radians. Agents can adjust the rudder only to the next closest two angles, preventing abrupt changes in direction.
@@
The episode ends when the maximum number of steps is reached.
Why is UTracking a difficult environment? UTracking poses a challenging benchmark due to partial observability, where agents can only detect targets within 450 m using noisy acoustic signals and share data through intermittent communication limited to 1500 m. The environment is highly stochastic, with sensor noise, communication dropouts, environmental perturbations, and unpredictable target motions. Effective tracking requires near-perfect coordination among agents, especially when following multiple fast-moving targets, as poor coordination can quickly lead to target loss. Finally, the long horizons of hour-scale missions (thousands of steps) demand learning policies that handle delayed consequences and remain robust to error accumulation in non-stationary multi-agent settings.
IV-C JaxLrauv UTracking